Ann, Ben and Cathy have their birthdays today. The sum of their ages is 23. The product of their ages is 113 more than the product of their ages on their birthday last year. What is the sum of the squares of their ages?
Found 2 solutions by ikleyn, mananth:
Answer by ikleyn(52786)

.
Ann, Ben and Cathy have their birthdays today. The sum of their ages is 23.
The product of their ages is 113 more than the product of their ages on their birthday last year.
What is the sum of the squares of their ages?
~~~~~~~~~~~~~~~


Simple calculations show that


    abc - (a-1)*(b-1)*(c-1) = ab + ac + bc - (a + b + c) + 1


Substituting here (a + b + c) = 23, we get this equality


    113 = ab + ac + bc - 23 + 1,

or

          ab + ac + bc = 113 + 23 - 1 = 135.


Now,  from the general formula


    (a + b + c)^2 = a^2 + b + c^2 + 2(ab + ac + bc)


we have


    a^2 + b^2 + c^2 = (a + b + c)^2 - 2(ab + ac + bc) = 23^2 - 2*135 = 259.


ANSWER.  The sum of squares of their ages is 259.

Solved.


////////////


@Manantrh successfully re-wrote my solution,  by adding nothing new to it . . . .


O,  no ---   . . .  by adding his name to it  (!)


Probably,  he (or she)  wanted to say that he  (or she)  was totally agree with my solution
and confirmed my calculations by repeating them again . . .



        Dear  @Mananth,  thank you for supporting and popularizing my solution  (!)




Answer by mananth(16946)

x+y+z=23
xyz=113+(x-1)(y-1)(z-1)
xyz = 113 +( xy-x-y+1)(z-1)
xyz= 113+ xyz-xy-xz+x-yz+y+z-1 (xyz) cancels off
xy+yz+zx =113+22
xy+yz+zx = 135
2(xy+yz+zx) =270
x^2+y^2+z^2=(x+y+z)^2-2(xy+yz+zx)
= 23^2-270 -259
the sum of the squares of their ages =259